Use the elimination method to find a general solution. x(t), y(t) for the given system. · = x + 2y dy = -4x - 3y dt
Answer #1

equation is in given system of differential ole - U+dy - -4*-zy -H) - (**) then given syorem can be written as ! (Driju - 2yX dla or to find so (8% + 2D + 5)% =0 - (ie they tady + 5y =o) yet), we nec to solve est ® auxiliary esuanion is not an tS to